OpenStudy (lexirebels):

Find the measure of one interior angle in a regular 53-gon.

OpenStudy (avi13):

let n sided polygon sum of intertior angles = 180(n-2) = 180(53-2)=9180 each interior angle = 180(n-2)/n=9180/53=173.2

OpenStudy (avi13):

sorry this angle is with centre to find interior angle consider it 2x and x +x+173.2 =180 (sum of angle of triangle) so x= 3.39 therefore interior angle is 2x= 6.79

OpenStudy (lexirebels):

Thanks